OpenStudy (anonymous):

A angle between line l and the line y-3=(-1/4)(x+2) measures 60 degrees. The line l contains the point (1,5). Find all possible equations of l.

OpenStudy (mertsj):

\[y-3=-\frac{1}{4}(x+2)\] \[y=-\frac{1}{4}x-\frac{1}{2}+3\] \[y=-\frac{1}{4}x+\frac{5}{2}\]

OpenStudy (anonymous):

One thing first. The method I've been taught is to use \[\tan \theta=\frac{ m1-m2 }{ 1+m1m2 }\] where m1 and m2 are slopes of intersecting lines.

OpenStudy (anonymous):

so I have  (Check if that's right, just in case.)